Emergency Help is what I need!!
3r-45=-13
4r+35=16
Solve the Elimination method
1. The solution is ____
2. There are infinitely many solution
3. There are no solution

Answer by Alan3354(69443)

Emergency Help is what I need!!
3r-45=-13
4r+35=16
Solve the Elimination method
1. The solution is ____
2. There are infinitely many solution
3. There are no solution
---------------------------
You have 2 equations in one variable, r.
Are these 2 separate problems?
-----------------
3r-45=-13
3r = 32
r = 32/3
--------------
4r+35=16
4r = -19
r = -19/4
---------------